# Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model

### v2

@@ Abstract
Then, $|\Psi_{A}(2)|$ becomes larger (resp. smaller) than $|\Psi_{A}(1)|$ when the nonlinear winding number is $\nu=1$ (resp. $\nu=0$), which indicates the localization (resp. delocalization) of the zero mode in a local sense. In the linear case ($b=0$), there is no $\Psi_{A}$ dependence in Eq. (19), and thus the winding number corresponds to the convergence to zero or the divergence. In nonlinear cases, the $\Psi_{A}$ dependence in Eq. (19) leads to the change of rates of amplification or attenuation of a zero mode, and thus both topological and trivial zero modes can converge to a fixed value. However, since the convergent value $\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$ must be larger (resp. smaller) than $|\Psi_{A}(1)|$ at $\nu=1$ (resp. $\nu=0$), the nonlinear winding number predicts the localization or delocalization of the zero mode in this parameter region, which is the bulk-edge correspondence of the nonlinear winding number. As clearly seen from Eq. (19), if we set the edge amplitude $|\Psi_{A}(1)|^{2}$ to be the critical amplitude of the nonlinear winding number, $|\Psi_{A}(1)|^{2}=w_{c}=-(a+d)/b$, $w_{c}$ also corresponds to the convergent value $w_{c}=\lim_{x\rightarrow\infty}|\Psi_{A}(x)|^{2}$ and we obtain a steady state where $|\Psi_{A}(x)|$ is independent of $x$. This is because the critical amplitude of the nonlinear winding number corresponds to the amplitude where the localization or delocalization of a zero mode is switched.
@@
Meanwhile, at $a>2$, the nonlinear effect destabilizes the convergence of a zero mode and induces the bifurcation to periodic solutions. This period-doubling bifurcation ubiquitously appears before a chaos transition. In fact, we find chaotic zero modes at $a\gtrsim 2.31$. Since the convergence values $\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$ are different from the bulk gap-closing points (the green curve in Fig. 2), one cannot predict such periodic and chaotic zero modes from the bulk topology, which implies the breakdown of the bulk-edge correspondence.
@@
We also analyze the behavior of zero modes by using cobweb plots, which visualize the nonlinear discrete dynamics in Eq. (19). Figures 3a,e show the edge modes in the parameter region where the winding number becomes nonzero in the linear limit $b\rightarrow 0$. In such a case, we obtain a localized zero mode fully decaying to $\lim_{x\rightarrow\infty}\Psi_{A}(x)=0$, which corresponds to a conventional topological edge mode. If we consider the case of $21<a<$ 2, the amplitude of the zero mode converges to a nonzero value as shown in Figs. 3b,f. Such a remaining amplitude at $x\rightarrow\infty$ is also reported in previous studies Hadad2016 ; Zhou2022 ; Sone2023 when the $w$-dependent nonlinear topological invariant signals the nonlinearity-induced topological phase transition. That is, the nonlinear winding number becomes zero (nonzero) at a smaller (larger) amplitude $w$ than the convergent value. While either amplitude gives a converging zero mode, the winding number $\nu(w)$ predicts its localization or anti-localization, which is the nonlinear bulk-edge correspondence. Considering larger $a$, we obtain periodic solutions shown in Figs. 3c,g. In particular, at $x\rightarrow\infty$, the zero mode described by the red polyline takes both larger and smaller values than the initial value, and thus one cannot judge the localization or delocalization of the zero mode. We also confirm the chaotic spatial distribution of the zero mode at $a=2.5$ (Figs. 3d,h). The spatial chaos of the zero mode is also characterized by the positive Lyapunov exponent $\lambda=0.728\ldots>0$ (see also Supplementary Materials). While we have focused on the semi-infinite system, the bulk-edge correspondence and its breakdown are also confirmed in the finite system of the nonlinear SSH model (see Supplementary Materials).
@@
Extension to long-range hoppings.— In the previous sections, we focus on the model only with nearest-neighbor hoppings and winding number $\nu(w)=1$ or 00. Long-range hoppings can lead to nonlinear winding numbers larger than one. In general linear systems, the absolute value of the winding number corresponds to the number of linearly independent edge modes Kane2014 . Meanwhile, since nonlinear systems have no superposition law, it has been unclear what corresponds to the absolute value of the nonlinear winding number. We here propose that it basically corresponds to the dimension of the stable manifold in the dynamical system describing a zero mode such as Eq. (19).
@@
To analyze the correspondence between the nonlinear winding number and the dimension of the stable manifold, we consider the extended nonlinear SSH model in Fig. 1. In this model, we add next-next-to-nearest-neighbor hoppings ($\alpha\Psi_{B}(x-2)$ to Eq. (Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model) and $\alpha\Psi_{A}(x+2)$ to Eq. (11)) to the nonlinear SSH model. Figure 4a presents the phase diagram of the extended SSH model in the linear limit $b=c=0$ Maffei2018 ; Perez2019 ; Hsu2020 . One can confirm that the winding number becomes $\nu=2$ at $\alpha>d-a$ and $\alpha>a$.
@@
which looks similar to the Duffing map Guckenheimer2013 and the Hénon map Henon2004 . We plot the vector field $(\Psi_{A}(x+1)-\Psi_{A}(x),\Psi_{A}(x+2)-\Psi_{A}(x+1))$ visualizing the deviations of the state variables at each step of this dynamical system in Figs. 4b-d. When the winding number is one in the linear limit, the fixed point at $\Psi_{A}=0$ is a saddle point, and thus its stable manifold is one-dimensional. In contrast, if the winding number is two in the linear limit, the fixed point at $\Psi_{A}=0$ is an attractor to which any points in its neighbor converge, and thus the dimension of its stable manifold is two. These results indicate the correspondence between the nonlinear winding number and the dimension of the stable manifold in weakly nonlinear cases. We note that in weakly nonlinear cases, the dimension of the stable manifold corresponds to the number of edge modes in the linear limit, and thus one can expect the correspondence between the winding number and the dimension of the stable manifold for further higher winding numbers.
Under stronger nonlinearity, we find the breakdown of the bulk-edge correspondence in the extended nonlinear SSH model in a sense similar to the bifurcation in the original nonlinear SSH model. If we consider $-(a+\alpha)<d<0$, $0<\alpha<a$, $b<0$, and $(a+\alpha)/|d|<2$, the model exhibits a nonlinearity-induced topological phase transition and a zero mode converging to $\Psi_{A}(x)=(a+\alpha+d)/|b|$. In this case, the nonlinear winding number becomes $\nu=1$ and the dimension of the stable manifold is also one as shown in Fig. 4d, which indicates the bulk-edge correspondence between the nonlinear winding number and the dimension of the stable manifold. However, at $(a+\alpha)/d>2$, the nonlinear winding number is unchanged, while the fixed point $\Psi_{A}(x)=(a+\alpha+d)/|b|$ is no longer a saddle point and becomes a fully unstable fixed point (see Supplementary Materials). At the critical parameter $(a+\alpha)/d=2$, the bifurcation to a periodic solution occurs. Therefore, the breakdown of the bulk-edge correspondence is induced by the bifurcation as in the original nonlinear SSH model.
@@
More general cases.— While we have focused on the nonlinear SSH model, the bulk-edge correspondence can be extended to more general nonlinear systems. Specifically, we here consider a sublattice-symmetric one-dimensional model whose nonlinear eigenvalue equation is described as
$E\left(\begin{array}[]{c}\bm{\psi}_{1}\\ \bm{\psi}_{2}\end{array}\right)=\left(\begin{array}[]{cc}0&A(f(\bm{\psi}))+D(f(\bm{\psi}))e^{-ik}\\ A(f(\bm{\psi}))+D(f(\bm{\psi}))e^{ik}&0\end{array}\right)\left(\begin{array}[]{c}\bm{\psi}_{1}\\ \bm{\psi}_{2}\end{array}\right),$ (28)
@@
where $\bm{\psi}_{1}$ and $\bm{\psi}_{2}$ are $N$-component vectors, and $A(f(\bm{\psi}))$ and $D(f(\bm{\psi}))$ are $N\times N$ matrices that are dependent on the nonlinear eigenvectors. We focus on the case that the dependence on the nonlinear eigenvector is determined only by a nonlinear function $f(\bm{\psi})$. Figure 5 shows the schematic of this lattice model, where each site belongs to either the first or second sublattice, and the interactions only exist between different sublattices. $A(f(\bm{\psi}))$ (resp. $D(f(\bm{\psi}))$) determines the amplitudes of intracell (resp. intercell) couplings. Under the right semi-infinite boundary condition, a nonlinear eigenvector with zero nonlinear eigenvalue follows
$\bm{\psi}_{1}(x+1)=T(f(\bm{\psi}_{1}(x)))\bm{\psi}_{1}(x),\ \ \bm{\psi}_{2}(x)=0,$ (29)
@@
with $T$ being the state-dependent transfer matrix, which is determined by $A(f(\bm{\psi}))$ and $D(f(\bm{\psi}))$ as
$T(f(\bm{\psi}_{1}(x)))=D^{-1}(f(\bm{\psi}_{1}(x)))A(f(\bm{\psi}_{1}(x))).$ (30)
@@
Thus, Eq. (29) only depends on the nonlinear function $f(\bm{\psi}_{1}(x))$ as in the nonlinear eigenvalue problem (Eq. (28)). Meanwhile, by assuming the Bloch ansatz $\bm{\psi}(x)=e^{ikx}\bm{\psi}(k)$ and fixing $f(\bm{\psi}(k))=w$ independently of the wavenumber $k$ (in the original nonlinear SSH model, we set $f(\bm{\psi}(k))=||\bm{\psi}(k)||^{2}$), one can obtain a nonlinear winding number $\nu(w)$ as in the generalized nonlinear SSH model. Then, if we fix the value of $f(\bm{\psi}_{A}(1))=w$ at the edge of the system, the dynamical system in Eq. (29) at $x=0$ becomes $\bm{\psi}_{A}(2)=T(w)\bm{\psi}_{A}(1),$ parametrized by $w$. As is inferred from the linear cases, the nonlinear winding number $\nu(w)$ corresponds to the number of eigenvalues of $T(w)$ whose absolute values are smaller than one (such correspondence can be shown from the argument principle, Supplementary Materials). Therefore, the correspondence between the nonlinear winding number $\nu(w)$ and amplification or attenuation around the edge site, $||\bm{\psi}(1)||<||\bm{\psi}(2)||$ or $||\bm{\psi}(1)||>||\bm{\psi}(2)||$, can be shown in more general nonlinear systems with arbitrary internal degrees of freedom. If we further assume that the tendency of the amplification or attenuation is unchanged in the limit of $x\rightarrow\infty$, one can show the bulk-edge correspondence in such a wide range of nonlinear systems. In contrast, such an assumption is broken by the chaos transition, which can also happen in this general case dependently on $T(w)$. Therefore, the breakdown of the bulk-boundary correspondence by the chaos transition also occurs in a wide range of nonlinear systems. We remain the mathematically rigorous argument as a future work.
@@
Discussion.— We revealed that nonlinear topological edge modes exhibit bifurcations to periodic solutions and chaos by analyzing the dynamical system describing the spatial distribution of zero modes. Such chaos transitions serve as the origin of the breakdown of the bulk-edge correspondence in nonlinear topology. We also proposed that the absolute value of a nonlinear topological invariant corresponds to the dimension of the stable manifold in a dynamical system describing the spatial distribution of zero modes, while such bulk-edge correspondence is also broken by the bifurcation.
@@
While we focused on one-dimensional systems, the analytical techniques used in this paper are applicable to arbitrary dimensions. Specifically, many higher-dimensional topological insulators are reduced to low-dimensional counterparts if we fix the wavenumber to a proper value; for example, the nonlinear Qi-Wu-Zhang model studied in previous research Sone2023 is equivalent to the nonlinear SSH model if we consider the wavenumber in the $y$ direction $k_{y}=0,\pi$. By using such reduction, one can extend the arguments of chaos in nonlinear topological edge modes to higher-dimensional nonlinear topological insulators. Therefore, the transition to chaos and the breakdown of the bulk-edge correspondence are ubiquitous in nonlinear systems. Meanwhile, our result shows that chaos, a well-known concept in nonlinear dynamical systems, can affect topological physics. Thus, we expect that interplays between nonlinear and topological physics can further uncover the characteristic behaviors in nonlinear topological insulators.
@@
It is noteworthy that $b\neq c$ leads to the difference in the critical amplitudes at which left and right edge modes appear (see Supplementary Materials). Such difference can correspond to a possible $\mathbb{Z}\times\mathbb{Z}$ classification, which is reminiscent of the topological classification of one-dimensional non-Hermitian systems Gong2018 ; Zhou2019 ; Kawabata2019 . However, since we have considered conserving nonlinear dynamics that correspond to Hermitian systems, the possible $\mathbb{Z}\times\mathbb{Z}$ classification is a genuinely nonlinear effect.
@@
There remains a possibility that one can define other topological invariants beyond the conventional wavenumber-space description and they can recover the bulk-edge correspondence in the chaotic region. Furthermore, while the models analyzed here are conservative dynamics where the sum of the amplitudes are unchanged in the time evolution, there are various dissipative (i.e., non-Hermitian-like) nonlinear systems in nature, such as biological fluids Marchetti2013 and oscillators Buck1988 ; Luo1991 . Therefore, the interplay between nonlinear and non-Hermitian topology Yuce2021 ; Zhu2022 ; Ezawa2022 and the extension of the bulk-edge correspondence to dissipative systems remain intriguing future issues.
@@
Therefore, we adapt this equation as the definition of the sublattice symmetry. We can also define the sublattice symmetries in the wavenumber-space description as
@@
where $P$ is a map representing the mirror operation. The mirror symmetry guarantees that every nonlinear eigenvector appears with its parity-inversion counterpart or is inversion symmetric in itself.
In the nonlinear SSH model, one can find the sublattice and mirror symmetries by considering linear maps $S$ and $P$. By defining $S$ as $S(\bm{\psi}(k))=\sigma_{z}\bm{\psi}(k)$ with $\sigma_{z}$ being the $z$-component of the Pauli matrix, one confirms the sublattice symmetry from the following calculation:
@@
Since the nonlinear SSH model holds the sublattice symmetry, one can define the nonlinear winding number.
Similarly, if $b$ is equal to $c$, one can confirm the mirror symmetry of the nonlinear SSH model by defining the mirror operator as $P=\sigma_{x}$ ($x$-component of the Pauli matrix). In fact, we obtain
@@
We here derive the recurrence relation in Eq. (19) in the main text that describes the spatial distribution of zero modes in the nonlinear SSH model. In the right semi-infinite system, the real-space description of the nonlinear eigenequation becomes as follows:

### v3

@@ Abstract
Then, $|\Psi_{A}(2)|$ becomes larger (resp. smaller) than $|\Psi_{A}(1)|$ when the nonlinear winding number is $\nu=1$ (resp. $\nu=0$), which indicates the localization (resp. anti-localization) of the zero mode in a local sense. In the linear case ($b=0$), there is no $\Psi_{A}$ dependence in Eq. (19), and thus the winding number corresponds to the convergence to zero or the divergence. In nonlinear cases, the $\Psi_{A}$ dependence in Eq. (19) leads to the change of rates of amplification or attenuation of a zero mode, and thus both topological and trivial zero modes can converge to a fixed value. However, such converging solutions from larger (smaller) $|\Psi_{A}(1)|$ than the convergent value $\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$ still represent spatially localized (anti–localized) nonlinear eigenvectors. Since one can confirm that the convergent value $\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$ must be larger (resp. smaller) than $|\Psi_{A}(1)|$ at $\nu=1$ (resp. $\nu=0$), the nonlinear winding number predicts the localization or anti-localization of the zero mode in this parameter region, which is the bulk–edge correspondence of the nonlinear winding number. As clearly seen from Eq. (19), if we set the edge amplitude $|\Psi_{A}(1)|^{2}$ to be the critical amplitude of the nonlinear winding number, $|\Psi_{A}(1)|^{2}=w_{c}=-(a+d)/b$, $w_{c}$ also corresponds to the convergent value $w_{c}=\lim_{x\rightarrow\infty}|\Psi_{A}(x)|^{2}$ and we obtain a steady state where $|\Psi_{A}(x)|$ is independent of $x$. This is because the critical amplitude of the nonlinear winding number corresponds to the amplitude where the localization or anti–localization of a zero mode is switched.
@@
Meanwhile, at $a>2$, the nonlinear effect destabilizes the convergence of a zero mode and induces the bifurcation to periodic solutions. Such destabilization of zero modes also breaks their localization, which induces the breakdown of the bulk–boundary correspondence. This period–doubling bifurcation ubiquitously appears before a chaos transition. In fact, we find chaotic zero modes at $a\gtrsim 2.31$ by numerically confirming the positive Lyapunov exponent (see Supplementary Note 2). Since the convergence values $\lim_{x\rightarrow\infty}|\Psi_{A}(x)|$ are different from the bulk gap-closing points (the orange curve in Fig. 2), one cannot predict such periodic and chaotic zero modes from the bulk topology, which implies the breakdown of the bulk–edge correspondence. We note that while some previous studies Engelhardt2017 ; Chaunsali2021 ; Ma2021 have discussed the instability and bifurcation in temporal dynamics of nonlinear edge modes, such temporal instability is not regarded as the breakdown of the bulk–boundary correspondence from the viewpoint of nonlinear eigenvalue problems. In contrast, the spatial chaos observed in the present model breaks the localized properties of edge modes, which induces the breakdown of the bulk-boundary correspondence unique to nonlinear systems. Interestingly, however, the temporal instability occurs at the bifurcation point in Fig. 2 and thus may also be related to the spatial instability (see Supplementary Note 3). We also note that since the spatial chaos is only seen in edge modes, such chaos modes are not directly relevant to the statistics of the bulk spectrum.
@@
We also analyze the behavior of zero modes by using cobweb plots, which visualize the nonlinear discrete dynamics in Eq. (19). Figures 3a,e show the edge modes in the parameter region where the winding number becomes nonzero in the linear limit $b\rightarrow 0$. In such a case, we obtain a localized zero mode fully decaying to $\lim_{x\rightarrow\infty}\Psi_{A}(x)=0$, which corresponds to a conventional topological edge mode. If we consider the case of $21<a<$ 2, the amplitude of the zero mode converges to a nonzero value as shown in Figs. 3b,f. Such a remaining amplitude at $x\rightarrow\infty$ is also reported in previous studies Hadad2016 ; Zhou2022 ; Sone2023 when the $w$-dependent nonlinear topological invariant signals the nonlinearity–induced topological phase transition. That is, the nonlinear winding number becomes zero (nonzero) at a smaller (larger) amplitude $w$ than the convergent value. While either amplitude gives a converging zero mode, the winding number $\nu(w)$ predicts its localization or anti–localization, which is the nonlinear bulk–edge correspondence. Considering larger $a$, we obtain periodic solutions shown in Figs. 3c,g. In particular, at $x\rightarrow\infty$, the zero mode described by the red polyline takes both larger and smaller values than the initial value, and thus one cannot judge the localization or anti-localization of the zero mode. We also confirm the chaotic spatial distribution of the zero mode at $a=2.5$ (Figs. 3d,h). The spatial chaos of the zero mode is also characterized by the positive Lyapunov exponent $\lambda=0.728\ldots>0$ (see also Supplementary Note 2). While we have focused on the semi–infinite system, the bulk–edge correspondence and its breakdown are also confirmed in the finite system of the nonlinear SSH model (see Supplementary Note 4). It is also noteworthy that both the edge modes and the chaos transition are robust against disorders (see Supplementary Note 5).
We emphasize that the chaos transition of zero modes can be universally seen in a wide range of nonlinear models. At the same time, such a transition is absent in some special models. For example, the SSH model with nonlinearity in the intercell hopping Hadad2016 ; Tuloup2020 ; Zhou2022 and the continuum model Sone2023 exhibit no chaos transitions, for which the bulk–boundary correspondence has been confirmed under fairly strong nonlinearity (we show such wide–range stability of edge modes in a nonlinear topological mechanics Ma2023 in Supplementary Note 6). Thus, the breakdown of the bulk–edge correspondence by chaos transitions has revealed a nontrivial nonlinear effect on topological edge modes.
@@
Extension to long-range hoppings.— In the previous sections, we focus on the model only with nearest–neighbor hoppings and the winding number $\nu(w)=1$ or 00. Long–range hoppings can lead to nonlinear winding numbers larger than one. In general linear systems, the absolute value of the winding number corresponds to the number of linearly independent edge modes Kane2014 . Meanwhile, since nonlinear systems have no superposition law, it has been unclear what corresponds to the absolute value of the nonlinear winding number. We here propose that it basically corresponds to the dimension of the stable manifold in the dynamical system describing a zero mode such as Eq. (19). While previous research Ma2021 has also indicated that topological edge modes can be regarded as orbits on stable manifolds, we reveal its correspondence to nonlinear topological invariants.
@@
To analyze the correspondence between the nonlinear winding number and the dimension of the stable manifold, we consider the extended nonlinear SSH model in Fig. 1. In this model, we add next–next–to–nearest–neighbor hoppings ($\alpha\Psi_{B}(x-2)$ to Eq. (Transition from topological to chaos in the nonlinear Su-Schrieffer-Heeger model) and $\alpha\Psi_{A}(x+2)$ to Eq. (11)) to the nonlinear SSH model. Figure 4a presents the phase diagram of the extended SSH model in the linear limit $b=c=0$ Zhang2017 ; Maffei2018 ; Hsu2020 . One can confirm that the winding number becomes $\nu=2$ at $\alpha>d-a$ and $\alpha>a$.
@@
which looks similar to the Duffing map Guckenheimer2013 and the Hénon map Henon2004 . We plot the vector field $(\Psi_{A}(x+1)-\Psi_{A}(x),\Psi_{A}(x+2)-\Psi_{A}(x+1))$ visualizing the deviations of the state variables at each step of this dynamical system in Figs. 4b–d. When the winding number is one in the linear limit, the fixed point at $\Psi_{A}=0$ is a saddle point, and thus its stable manifold is one-dimensional. In contrast, if the winding number is two in the linear limit, the fixed point at $\Psi_{A}=0$ is an attractor to which any points in its neighbor converge, and thus the dimension of its stable manifold is two. These results indicate the correspondence between the nonlinear winding number and the dimension of the stable manifold in weakly nonlinear cases. We note that in weakly nonlinear cases, the dimension of the stable manifold corresponds to the number of edge modes in the linear limit, and thus one can expect the correspondence between the winding number and the dimension of the stable manifold for further higher winding numbers.
Under stronger nonlinearity, we find the breakdown of the bulk–edge correspondence in the extended nonlinear SSH model in a sense similar to the bifurcation in the original nonlinear SSH model. If we consider $-(a+\alpha)<d<0$, $0<\alpha<a$, $b<0$, and $(a+\alpha)/|d|<2$, the model exhibits a nonlinearity–induced topological phase transition and a zero mode converging to $\Psi_{A}(x)=(a+\alpha+d)/|b|$. In this case, the nonlinear winding number becomes $\nu=1$ and the dimension of the stable manifold is also one as shown in Fig. 4d, which indicates the bulk-edge correspondence between the nonlinear winding number and the dimension of the stable manifold. However, at $(a+\alpha)/d>2$, the nonlinear winding number is unchanged, while the fixed point $\Psi_{A}(x)=(a+\alpha+d)/|b|$ is no longer a saddle point and becomes a fully unstable fixed point (see Supplementary Note 7). At the critical parameter $(a+\alpha)/d=2$, the bifurcation to a periodic solution occurs. Therefore, the breakdown of the bulk–edge correspondence is induced by the bifurcation as in the original nonlinear SSH model. One can also find that the long–range hopping can increase the number of anti-localized modes corresponding to the higher dimension of the stable manifold (see Supplementary Note 8).
@@
More general cases.— While we have focused on the nonlinear SSH model, the bulk–edge correspondence can be extended to more general nonlinear systems. Specifically, we here consider a sublattice–symmetric one-dimensional model whose nonlinear eigenvalue equation is described as
$E\left(\begin{array}[]{c}\bm{\psi}_{1}\\ \bm{\psi}_{2}\end{array}\right)=\left(\begin{array}[]{cc}0&A^{\dagger}(g(\bm{\psi}))+D^{\dagger}(g(\bm{\psi}))e^{-ik}\\ A(g(\bm{\psi}))+D(g(\bm{\psi}))e^{ik}&0\end{array}\right)\left(\begin{array}[]{c}\bm{\psi}_{1}\\ \bm{\psi}_{2}\end{array}\right),$ (28)
@@
where $\bm{\psi}_{1}$ and $\bm{\psi}_{2}$ are $N$–component vectors, and $A(g(\bm{\psi}))$ and $D(g(\bm{\psi}))$ are $N\times N$ matrices that are dependent on the nonlinear eigenvectors. We focus on the case that the dependence on the nonlinear eigenvector is determined only by a nonlinear function $g(\bm{\psi})$. Figure 5 shows the schematic of this lattice model, where each site belongs to either the first or second sublattice, and the interactions only exist between different sublattices. $A(g(\bm{\psi}))$ (resp. $D(g(\bm{\psi}))$) determines the amplitudes of intracell (resp. intercell) couplings. Under the right semi–infinite boundary condition, a nonlinear eigenvector with zero nonlinear eigenvalue follows
$\bm{\psi}_{1}(x+1)=T(g(\bm{\psi}_{1}(x)))\bm{\psi}_{1}(x),\ \ \bm{\psi}_{2}(x)=0,$ (29)
@@
with $T$ being the state–dependent transfer matrix, which is determined by $A(g(\bm{\psi}))$ and $D(g(\bm{\psi}))$ as
$T(g(\bm{\psi}_{1}(x)))=D^{-1}(g(\bm{\psi}_{1}(x)))A(g(\bm{\psi}_{1}(x))).$ (30)
@@
Thus, Eq. (29) only depends on the nonlinear function $g(\bm{\psi}_{1}(x))$ as in the nonlinear eigenvalue problem (Eq. (28)). Meanwhile, by assuming the Bloch ansatz $\bm{\psi}(x)=e^{ikx}\bm{\psi}(k)$ and fixing $g(\bm{\psi}(k))=w$ independently of the wavenumber $k$ (in the original nonlinear SSH model, we set $g(\bm{\psi}(k))=||\bm{\psi}(k)||^{2}$), one can obtain a nonlinear winding number $\nu(w)$, which also indicates the possible usefulness of fixing nonconservative quantity. Then, if we fix the value of $g(\bm{\psi}_{A}(1))=w$ at the edge of the system, the dynamical system in Eq. (29) at $x=0$ becomes $\bm{\psi}_{A}(2)=T(w)\bm{\psi}_{A}(1)$ parametrized by $w$. As is inferred from the linear cases, the nonlinear winding number $\nu(w)$ corresponds to the number of eigenvalues of $T(w)$ whose absolute values are smaller than one (such correspondence can be shown from the argument principle, Supplementary Note 9). Therefore, the correspondence between the nonlinear winding number $\nu(w)$ and amplification or attenuation around the edge site, $||\bm{\psi}(1)||<||\bm{\psi}(2)||$ or $||\bm{\psi}(1)||>||\bm{\psi}(2)||$, can be shown in more general nonlinear systems with arbitrary internal degrees of freedom. If we further assume that the tendency of the amplification or attenuation is unchanged in the limit of $x\rightarrow\infty$, one can show the bulk–edge correspondence in such a wide range of nonlinear systems. In contrast, such an assumption is broken by the chaos transition, which can also happen in this general case dependently on $T(w)$. Therefore, the breakdown of the bulk–boundary correspondence by the chaos transition also occurs in a wide range of nonlinear systems. We remain the mathematically rigorous argument as a future work.
While we have focused on sublattice-symmetric systems, topological phases and edge modes can be realized under other symmetries, such as the time–reversal and spatial–inversion symmetries. Under those symmetries, one can use the nonlinear Berry phase as a topological invariant Zhou2022 instead of the nonlinear winding number (Eq. (9)). Even without the sublattice symmetry, one can still consider spatial dynamics of edge modes similar to Eq. (19). In such spatial dynamics, it is expected that one can also observe chaos transitions, and thus the breakdown of the bulk–boundary correspondence by chaos transitions universally occurs independently of the symmetries of nonlinear systems.
We also note that there are various nonlinear systems without the $U(1)$ symmetry, such as fluids Marchetti2013 , nonlinear oscillators Luo1991 , mechanical lattices Ma2023 , and electrical circuits Zhou2022 . Such broken $U(1)$ symmetry may alter the topological classification of nonlinear systems, as is the case for interacting bosonic cases. Furthermore, a previous paper Sone2023 has shown that the $U(1)$ symmetry is necessary to obtain the wavenumber–space description of the nonlinear eigenvalue problem (cf. Eq. (8)) from the Bloch ansatz. Possible extension of the topological invariants to one-dimensional systems without $U(1)$ symmetry has been discussed in some previous studies Zhou2022 ; Zhou2024 by assuming modulated waves instead of the Bloch ansatz. By using a similar technique, we expect that our results of the present paper can be extended to more various systems even without the $U(1)$ symmetry.
@@
Discussion.— We revealed that nonlinear topological edge modes exhibit bifurcations to periodic solutions and chaos by analyzing the dynamical system describing the spatial distribution of zero modes. Such chaos transitions serve as the origin of the breakdown of the bulk–edge correspondence in nonlinear topology. We also proposed that the absolute value of a nonlinear topological invariant corresponds to the dimension of the stable manifold in a dynamical system describing the spatial distribution of zero modes, while such bulk–edge correspondence is also broken by the bifurcation.
@@
While we focused on one–dimensional systems, the analytical techniques used in this paper are applicable to arbitrary dimensions. Specifically, many higher–dimensional topological insulators are reduced to low–dimensional counterparts if we fix the wavenumber to a proper value; for example, the nonlinear Qi–Wu–Zhang model studied in previous research Sone2023 is equivalent to the nonlinear SSH model if we consider the wavenumber in the $y$ direction $k_{y}=0,\pi$ (see Methods). By using such reduction, one can extend the arguments of chaos in nonlinear topological edge modes to higher–dimensional nonlinear topological insulators. Furthermore, even at $k_{y}\neq 0,\pi$, one can obtain a nonlinear dynamical system in the spatial direction similar to Eq. (19) (see also Methods). Therefore, the transition to chaos and the breakdown of the bulk–edge correspondence are ubiquitous in nonlinear systems. Meanwhile, our result shows that chaos, a well–known concept in nonlinear dynamical systems, can affect topological physics. Thus, we expect that interplays between nonlinear and topological physics can further uncover the characteristic behaviors in nonlinear topological insulators.
@@
It is noteworthy that $b\neq c$ leads to the difference in the critical amplitudes at which left and right edge modes appear (see Supplementary Notes 10 and 11). Such difference can correspond to a possible $\mathbb{Z}\times\mathbb{Z}$ classification, which is reminiscent of the topological classification of one-dimensional non-Hermitian systems Gong2018 ; Zhou2019 ; Kawabata2019 . However, since we have considered conserving nonlinear dynamics that correspond to Hermitian systems, the possible $\mathbb{Z}\times\mathbb{Z}$ classification is a genuinely nonlinear effect.
@@
There remains a possibility that one can define other topological invariants beyond the conventional wavenumber–space description and they can recover the bulk–edge correspondence in the chaotic region. Furthermore, while the models analyzed here are conservative dynamics where the sum of the amplitudes are unchanged in the time evolution, there are various dissipative (i.e., non–Hermitian–like) nonlinear systems in nature, such as biological fluids Marchetti2013 and oscillators Luo1991 . Therefore, the interplay between nonlinear and non–Hermitian topology Yuce2021 ; Zhu2022 ; Ezawa2022 and the extension of the bulk-edge correspondence to dissipative systems remain intriguing future issues.
@@
Therefore, we adapt this equation as the definition of the sublattice symmetry. We can also define the sublattice symmetries in the wavenumber–space description as
@@
where $P$ is a map representing the mirror operation. The mirror symmetry guarantees that every nonlinear eigenvector appears with its parity–inversion counterpart or is inversion symmetric in itself.
In the nonlinear SSH model, one can find the sublattice and mirror symmetries by considering linear maps $S$ and $P$. Here we denote the right–hand sides of Eq. (18) by ${\bf f}(k;(\psi_{A},\psi_{B}))$. By defining $S$ as $S(\bm{\psi}(k))=\sigma_{z}\bm{\psi}(k)$ with $\sigma_{z}$ being the $z$–component of the Pauli matrix, one confirms the sublattice symmetry from the following calculation:
@@
Since the nonlinear SSH model holds the sublattice symmetry, one can define the nonlinear winding number. The sublattice symmetry of the nonlinear SSH model also justifies the assumption that the nonlinear eigenvalue of an edge mode is zero.
Similarly, if $b$ is equal to $c$, one can confirm the mirror symmetry of the nonlinear SSH model by defining the mirror operator as $P=\sigma_{x}$ ($x$–component of the Pauli matrix). In fact, we obtain
@@
This mirror symmetry is related to the appearance of the gapless point at $k=0$ or $k=\pi$.
@@
We here derive the recurrence relation in Eq. (19) in the main text that describes the spatial distribution of zero modes in the nonlinear SSH model. In the right semi–infinite system, the real-space description of the nonlinear eigenequation becomes as follows: